OpenStudy (anonymous):

The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?

OpenStudy (ivettef365):

on the word you have 11 letters in total 4 letters are (i)

OpenStudy (ivettef365):

probability to draw an i would be 4/11, then probability to draw a second (i) 3/10, because now there are 10 ltrs left and 3 (i) left

OpenStudy (kropot72):

There are four i's in the word and the word has 11 letters. The probability of drawing an i on the first draw is 4/11. Given that an i was drawn on the first draw, the probability of drawing an i on the second draw is 3/10. The required probability is found from the product of the 2 values of probability: \[P(2\ letter\ i)=\frac{4}{11} \times \frac{3}{10}\]

OpenStudy (ivettef365):

so multiply 4/11 x 3/10

OpenStudy (anonymous):

6/55 ?

OpenStudy (kropot72):

You answer is correct :)

OpenStudy (anonymous):

thanks so much! both of you!

OpenStudy (kropot72):

You're welcome :)